# Run-Time Monitoring of ERTMS/ETCS Control Flow by Process Mining

### v1

@@ 5.2 Results and discussion
We start the analysis of results from Table 1, which presents accuracy and v-measure scores for different clustering algorithms under windows sizes of 5, 10, 15 (see Section 4.2). For the sake of brevity, the table does not report results for all clustering algorithms. DBSCAN and MiniBatchKMeans are omitted as (in our experiments) they are strictly worse than HDBSCAN and K-Means, respectively. Optics, MeanShift and AffinityPropagation are not appearing as well as outperformed by algorithms in the table. Overall, accuracy and V-measure scores are consistent: aside from a few corner cases, the higher the accuracy, the higher the V-measure. Additionally, relying on larger window sizes allows for a noticeable increase in metric scores (i.e., the user can compare the 1st against the 5th row of the table for accuracy, or the 2nd against the 6th row for v-measure), regardless of the algorithm. The spectral clustering achieves the highest accuracy and V-measure scores with 50 clusters, as can be seen by the bolded quantities in the table. The last remark regards the number of clusters to be used to train algorithms: overall, the more clusters, the better, as the algorithm has more freedom to create even small clusters that have a very good homogeneity i.e., contain only data points that describe very specific features of a class.
@@
Another view on results can be discussed with the aid of Figure 5(a), Figure 5(b), Figure 5(c), which focus on the behavior of the Spectral Clustering with 50 clusters (the best clustering strategy from the previous discussion) under different window sizes and under the injection of different faults. The three figures do not show accuracy nor v-measure scores: instead, they show the relationship of:
  • $S_{ARBC}$, $S_{EVC}$, $S_{HRBC}$, $S_{RTM}$, averaged over data points in test sets, painted as blue-striped, orange solid, green diagonally-striped, and with a purple gradient, respectively;
  • experiments with the injection of arbc (first batch of bars in plots), evc (second batch), hrbc (third batch, rtm (last batch) faults
The reader may assume that the activation of a fault of a specific type $ft\in\{ARBC,EVC,HRBC,RTM\}$ will lead to an high value of $S_{ft}$, leaving other metrics behind, or even at zero. Conversely, the bar charts show that in some configurations this is not true. It is the case of the sets of bars arbc fault in Figure 5(a), for which the bar corresponding to $S_{ARBC}$ is not the highest, tied with the $S_{HRBC}$ bar and shorter than the $S_{RTM}$ bar. A similar behavior can be observed in the same figure with and evc fault bars, where the $S_{EVC}$ and $S_{RTM}$ quantities are on par. Overall, bar charts in Figure 5(a), Figure 5(b), and Figure 5(c) show us that non-comnformances, or misalignments, are happening more frequently in the faulty component, and in those that follow the faulty component in the faulty trace. When injecting an EVC fault (second set of bars in all figures), there is just a small likelihood of having misalignments in the ARBC component (which comes before the EVC component in the trace), but there is a consistent portion of misalignments that are happening also in the HRBC and RTM components, which are exercised after the faulty component. While in some cases these cascading effects of the faut activation outside the faulty component are marginal, there may be cases - as discussed before in Figure 5(a) - in which these cascading effects trigger more misalignments outside the faulty component than inside it.
In conclusion, our experimentation outlined that:
@@
  • Online conformance checking of the online traces with the Petri net in Fig. 4 allowed us to extract diagnoses that can be accurately handled by our clustering process to predict the fault that occurred in the system, achieving up to 94% accuracy and 81% V-measure;
@@
  • Our explanation process allowed us to identify the misbehaving components of different online traces. The larger the window size, the more accurate the localization is with respect to the specific faulty component.
These results open the opportunity to experiment with more process discovery and conformance checking algorithms to evaluate their impact on the subsequent steps of the methodology.
@@ 6 Conclusions
In this paper, we presented a novel approach for run-time control-flow anomaly detection in ERTMS/ETCS L2 using process mining. By leveraging execution traces and conformance checking, the behavior of the system can be monitored to identify deviations from expected operational models. Our method also integrates unsupervised machine learning to support anomaly localization, offering insights into the root causes of observed deviations with a high degree of accuracy and explainability. The RBC/RBC Handover scenario has been chosen as a good fitting use-case of complex coordination tasks in train control; as such, it served as a reference case-study to validate our approach through a proof-of-concept, demonstrating its feasibility and effectiveness in realistic settings.
This work contributes to enhancing the dependability and fault-tolerance of safety-critical railway systems by bridging the gap between offline verification and online assurance. In line with the growing need for more resilient and intelligent railway software, our future research will focus on incorporating self-adaptive capabilities, wherein detected anomalies may trigger autonomous mitigation actions or adaptation strategies. Furthermore, the integration of our monitoring framework within digital twin architectures of ERTMS/ETCS subsystems is foreseen, enabling predictive diagnostics and closed-loop assurance [12, 34].
@@ 6.0.1 Acknowledgements
The work of Francesco Vitale and Nicola Mazzocca was partly supported by the Spoke 9 “Digital Society & Smart Cities” of ICSC - Centro Nazionale di Ricerca in High Performance-Computing, Big Data and Quantum Computing, funded by the European Union - NextGenerationEU (PNRR-HPC, CUP: E63C22000980007). The work of Francesco Flammini was partly supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contracts no. 23.00321 (Academics4Rail) and 24.00528 (PhDs EU-Rail); those projects have been selected within the European Union’s Horizon Europe research and innovation programme under grant agreements no. 101121842 and 101175856, respectively. The work of Tommaso Zoppi was partly supported by the the project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union – NextGenerationEU, and by the Cognitive Safety with Point Clouds (CogniSafe3D) Eurostars 3, Call 6 by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the funding agencies, which cannot be held responsible for them.

### v2

@@ 5.2 Anomaly Detection Results
Table 1 reports the anomaly detection results in terms of accuracy and V-measure for different window sizes and clustering algorithms. We limit our analysis to the best-performing clustering algorithms requiring the manual tuning of the number of clusters (K-Means, WARD, Spectral Clustering, and BIRCH) and the best-performing density-based clustering algorithm (HDBScan).
First, let us consider the impact of increasing the number of clusters in K-Means, WARD, Spectral Clustering and BIRCH on the evaluation metrics. Except for BIRCH, increasing the number of clusters leads to better performance, with Spectral Clustering peaking at 0.94 accuracy and 0.81 V-measure with 50 clusters. Such high performance is due to the ability of the algorithm to build specific, homogeneous, and complete clusters. However, it is worth noting that a higher cluster number increases the computational complexity and may lead to overfitting due to the specificity of the clusters being built. Therefore, selecting a lower number of clusters could be safer in terms of generalization capability and speed while maintaining a reasonably high accuracy, as Spectral Clustering achieves 0.82 accuracy and 0.61 V-measure with only 10 clusters.
Second, let us discuss the impact of the window size on the evaluation metrics. On the one hand, a window size equal to 5 leads to worse performance, with the lowest accuracy and V-measure being 0.61 and 0.32 for K-Means. On the other hand, increasing the window size leads to higher accuracy and V-measure across all the clustering algorithms — except for HDBScan, whose accuracy and V-measure, respectively, drop from 0.69 to 0.34 for window size equal to 5 and 0.37 to 0.09 for window size equal to 15. Hence, while K-Means, WARD, Spectral Clustering, and BIRCH are able to build better clusters through the additional conformance checking diagnoses yielded by a larger window size, HDBScan is unable to effectively use such information. In conclusion, it is worth noting that, despite the majority of algorithms achieving better accuracy and V-measure, increasing the window size generally leads to higher conformance checking times.
@@
In conclusion, online conformance checking of the online traces with the Petri net in Fig. 4 allowed us to extract diagnoses that can be accurately handled by our clustering process to predict the fault that occurred in the system, achieving up to 0.94 accuracy and 0.81 V-measure.
5.3 Anomaly Explanation Results
The bar plots of Figs. 5(a), 5(b), and 5(c) show the anomaly explanations $S_{ARBC}$, $S_{EVC}$, $S_{HRBC}$ and $S_{RTM}$ for window sizes 5, 10, and 15, respectively. These results are based on the clusters built with Spectral Clustering using 50 clusters (the best clustering strategy from the previous discussion). $S_{ARBC}$, $S_{EVC}$, $S_{HRBC}$, and $S_{RTM}$ are divided into four batches. Each batch collects the anomaly explanations related to the test traces of a specific fault type. For example, the first batch on the left-hand side of each bar plot collects the anomaly explanations of the $N_{anom,ARBC}$ traces.
Although it is expected that the activation of a fault within a specific component compcomp leads to high $S_{comp}$ with a clear separation from the other explanations, some fault type-window size configurations fail to achieve this result. For example, the results of a window size equal to 5 shown in Fig. 5(a) show that $S_{ARBC}$ is lower than $S_{RTM}$ despite the activation of an ARBC fault. However, this does not occur for window sizes equal to 10 (Fig. 5(b)) and 15 (Fig. 5(c)), which is consistent with the accuracy and V-measure reported in Table 1. Still, while window sizes equal to 10 and 15 allow better localization of the ARBC and EVC faults, the separation of the different anomaly explanations is not so clear. On the other hand, the HRBC and RTM faults always lead to, respectively, high $S_{HRBC}$ and $S_{RTM}$, and, in both cases, a very clear separation of anomaly explanations regardless of the window size. This suggests that faults in specific components are more likely to propagate to other components, making the labeling and localization of the fault more challenging in any window size scenario.
@@ 5.3 Anomaly Explanation Results
In summary, our explanation process allowed us to identify the misbehaving components of different online traces. The larger the window size, the more accurate the localization is with respect to the specific faulty component.
@@ 6 Conclusions
In this paper, we presented a novel approach for run-time control-flow anomaly detection in ERTMS/ETCS L2 using process mining. By leveraging execution traces and conformance checking, the behavior of the system can be monitored to identify deviations from expected execution models. Our method also integrates unsupervised machine learning to support anomaly detection and localization, offering insights into the root causes of observed deviations with a high degree of accuracy and explainability. The RBC/RBC Handover scenario has been chosen as a good-fitting use case of complex coordination tasks in train control; as such, it served as a reference case-study to validate our approach through a proof-of-concept, demonstrating its feasibility and effectiveness in realistic settings.
This work contributes to enhancing the dependability and fault-tolerance of safety-critical railway systems by bridging the gap between offline verification and online assurance. In line with the growing need for more resilient and intelligent railway software, our future research will focus on incorporating self-adaptive capabilities, wherein detected anomalies may trigger autonomous mitigation actions or adaptation strategies. Furthermore, the integration of our monitoring framework within digital twin architectures of ERTMS/ETCS subsystems is foreseen, enabling predictive diagnostics and closed-loop assurance [debenedictis2023dtadiiot, vitale2025pmdt].
@@ 6.0.1 Acknowledgements
The work of Francesco Vitale and Nicola Mazzocca was partly supported by the Spoke 9 “Digital Society & Smart Cities” of ICSC - Centro Nazionale di Ricerca in High Performance-Computing, Big Data and Quantum Computing, funded by the European Union - NextGenerationEU (PNRR-HPC, CUP: E63C22000980007). The work of Francesco Flammini was partly supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contracts no. 23.00321 (Academics4Rail) and 24.00528 (PhDs EU-Rail); those projects have been selected within the European Union’s Horizon Europe research and innovation programme under grant agreements no. 101121842 and 101175856, respectively. The work of Tommaso Zoppi was partly supported by the project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union – NextGenerationEU, and by the Cognitive Safety with Point Clouds (CogniSafe3D) Eurostars 3, Call 6 by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the funding agencies, which cannot be held responsible for them.